OpenStudy (anonymous):

a boat moves from p to q (downstream) in 2.5 h and from q to p (upstream) in 5 h respectively when the engie is producing maximum speed on the boat in each case . if due to increased water flow the speed of flow of the river gets doubled how long will it take the boat(moving with possible speed) to move from (a) p to q (downstream) (b) q to p (upstream)

OpenStudy (shubhamsrg):

let distance be d..let boat speed be v_b ,,and river speed be v_r now for upstream v_b - v_r = d/5 and for downstream v_b + v_r = d/2.5 = 2d/5 on solving , v_b = 3d/10 and v_r = d/10 now new speed of river,v_R = 2d/10 =d/5 let unknown times be t_d and t_u again,,for downstream,, v_b + v_R = d /t_d => substitute values and find t_d and for upstream, v_b - v_R = d/t_u => subst. values and find t_u

OpenStudy (anonymous):

it is 3d/5 is it??

OpenStudy (shubhamsrg):

what is 3d/5 ?

OpenStudy (anonymous):

2d/5+d/5=3d/5

OpenStudy (shubhamsrg):

well (2v_b)= 3d/5 ..

OpenStudy (anonymous):

wat? am i correct

OpenStudy (shubhamsrg):

am not getting what you're asking ?

OpenStudy (anonymous):

v_b - v_r = d/5 and for downstream v_b + v_r = d/2.5 = 2d/5 on solving , v_b = 3d/10 and v_r = d/10

OpenStudy (anonymous):

v_b=3/5 is it?

OpenStudy (shubhamsrg):

d is not given,,how can you say 3d/10 = 3/5

OpenStudy (anonymous):

sorry it 3d/5

OpenStudy (shubhamsrg):

its 3d/10 !! how in this world you come up with 3d/5